OpenStudy (anonymous):

sin^2beta/1+cosbeta =1

OpenStudy (anonymous):

\[\frac{ \sin ^2\beta }{ 1+\cos \beta }=\frac{ 1-\cos ^2 \beta }{ 1+\cos \beta }=\frac{ \left( 1+\cos \beta \right)\left( 1-\cos \beta \right) }{ 1+\cos \beta} =?=1\] find beta.